OpenStudy (cherie_magee):

Given: 3x + y = 1. Solve for y. y = -3x + 1 y = -3x - 1 y = 3x - 1

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (cherie_magee):

weeelll i kinda think its a but im not sure @skullpatrol

OpenStudy (skullpatrol):

You are right :)

OpenStudy (cherie_magee):

are you sure because im not @skullpatrol

OpenStudy (skullpatrol):

Yes, I am sure because if you add -3x to both sides of the equation then you will get your answer.

OpenStudy (skullpatrol):

Given: $$3x + y = 1$$ Solve for y. $$-3x+3x+y=-3x+1$$ $$0 +y=-3x+1$$ $$y=-3x+1$$

OpenStudy (skullpatrol):

This is known as the addition property of equality; where you are allowed to add the same real number to both sides of the equation.

OpenStudy (skullpatrol):

If a,b and c are any real numbers and a = b, then a + c = b +c and c + a = c + b. If the same number is added to equal numbers the sums are equal.